OpenStudy (anonymous):

-4[(-3)(-1 )- (-1)^6(-4)] /(4-10)^2

OpenStudy (anonymous):

Hi @alyygirl. Hope you have been well. Let's work through this one? Or are you looking for just an answer...?

OpenStudy (anonymous):

lets work through

OpenStudy (anonymous):

-4 [(-3)(-1) - -1^6(-4)] / (4-10)^2 Let's split it up into manageable sections so we can do it step by step. I didn't change anything, just added a bunch of spaces to separate the parts, can you see that so far?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

Ok so we can simplify the (-3)(-1) to be 3 right? -4 [ 3 - -1^6(-4)] / (4-10)^2

OpenStudy (anonymous):

-4 [(-3)(-1) - -1^6(-4)] / (4-10)^2 (3)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

-4 [ 3 - -1^6(-4)] / (4-10)^2 -1^6 1(-4) -4 -4 [ 3 - -4 / (4-10)^2

OpenStudy (anonymous):

Simplifying the -1^6(-4) in the last part I showed.

OpenStudy (anonymous):

which is -4

OpenStudy (anonymous):

-4 [3 - (-4)] / (4-10)^2 - * - + 3 + 4 7 -4 [ 7 ] / (4-10)^2 I really hope this is making sense. Do you have any questions about how I came to -4(7)/(4-10)^2 so far?

OpenStudy (anonymous):

no im good

OpenStudy (anonymous):

That's awesome!! I felt like I was doing a horrid job haha

OpenStudy (anonymous):

nope but the way its written is confusing lol

OpenStudy (anonymous):

(4-10)^2 = -6^2 = 36 -4(7)/36 -28/36 -14/18 -7/9 Our answer should be -7/9 but I'm pretty sure I made a mistake somewhere...

OpenStudy (anonymous):

ok so it is -7/9 or your not sure?

OpenStudy (anonymous):

I am not 100% sure. More like 25%. Let me do it on paper real quick ok? =0P

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

Yep that's right!

OpenStudy (anonymous):

I'm good.

OpenStudy (anonymous):

100% sure.

OpenStudy (anonymous):

o thanks :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

My pleasure.

OpenStudy (anonymous):

another question?

OpenStudy (anonymous):

actually check my work

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Evaluate the following expression when a = -3, b = 1, c = 3: -3bc over bc^2 - c

OpenStudy (anonymous):

-39/166?

OpenStudy (anonymous):

\[-3bc/bc ^{2} - c\] Right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[-3(1)(3)/(1)(3^{2}) - 3\] Where is the a in this equation lol?

OpenStudy (anonymous):

expression, not equation

OpenStudy (anonymous):

its the -3 i guess

OpenStudy (anonymous):

idk weird

OpenStudy (anonymous):

9/6 or 3/2 is the reduced form of that expression.

OpenStudy (anonymous):

-9/6 or -3/2 sorry.

OpenStudy (anonymous):

ok